At first there were 17 boring beetles. The number of boring beetles doubled every day.
(a) Write an equation that describes the number of beetles there would be after n days.
(b) Using the equation in (a), determine the number of beetles there would be after 30 days.

Answer by math_tutor2020(3817)


Answers:
(a) P = 17*2^n
(b) 18,253,611,008 (roughly 18.25 billion)

=======================================================================
Explanation for Part (a)

One possible exponential equation template is P = a*b^n
where,
n = number of days
P = number of beetles
a = starting amount
b = growth rate when b > 1

In this case,
a = 17
b = 2, meaning the population doubles each time n goes up by 1.

We go from P = a*b^n to P = 17*2^n

If n = 0, then P = 17
If n = 1, then P = 34
If n = 2, then P = 68
and so on.

====================================
Explanation for Part (b)

Plug n = 30 into the equation we just found.
P = 17*2^n
P = 17*2^30
P = 18,253,611,008
This population count is roughly 18.25 billion.

I don't know much about beetles, or insects, to know if this population count is realistic or not.